Style Specifications for Electronic Submissions to the New York Journal of Mathematics

Mark Steinberger

Abstract. We give style specifications for submission of articles to the New York Journal of Mathematics.

We give style specifications for submission of articles in either AMS-L^{A}TEX or AMS-TEX. We assume a basic familiarity with writing mathematics in whichever of the two formats you choose to use. The fine points of writing in AMS-TEX are discussed in [Spi90]. For AMS-L^{A}TEX, one should read [Lam86] and [AUG90], and perhaps take a look at [Spi90], as well.

Contents

1. General Rules 1

2. Theorem-like Environments 3 2.1. AMS-L^{A}TEX 2.2. AMS-TEX

3. Commutative Diagrams, Tables, and Graphics 4

4. Bibliographic References 5

References 6

1. General Rules

• Your paper should be sent in a single file, including all author-written macros.^{1} No macro packages may be input other than those explicitly listed here. Copies of these packages will be available from our archive.

1Graphical input of certain types may be submitted in separate files. Please see Section 3, below.

• Papers may be submitted in either AMS-TEX or AMS-L^{A}TEX. In either case, the paper should be prepared using one of the style files provided in our archive. For AMS-L^{A}TEX papers, use the primary documentstyle nyj.sty, together with the subsidiary file nyj10.sty. For AMS-TEX papers, use nyjamst.sty. The nyj and nyjamst styles are based on the amsart and amsppt styles created by the A.M.S. But the differences are sufficient that papers should be prepared using our styles, rather than the A.M.S. ones.

• You should also use the template for topmatter provided in our archive for the flavor of TEX you have chosen.

• All author-written commands should be via \newcommand in AMS-L^{A}TEX or \define in AMS-TEX. They must not conflict with command definitions in the format or document style: \renewcommand, \redefine, and the various forms of TEX’s \def should not be used.

• All author-written commands should appear immediately after the topmatter, as indicated in the templates for topmatter of submitted papers. Command definitions should not appear in the body of the paper, at the heads of sections, etc.^{2}

• No formatting commands (such as forced page breaks or line breaks, forced vertical or horizontal skips, etc.) are permitted. Since slightly different versions of the macros generating TEX’s formats will give different line and page breaks, the fine tuning by the author should be limited to hyphenation commands, suggested breaks, rewordings, and the placement of long mathematical expressions in displays.

• Capitalization in the title of the paper itself and in the titles of parts and sections should follow the convention that the first letter in each major word is upper case and the remaining letters are lower case. The same convention should be used for user-specified running heads. Similarly, authors’ names should be entered in caps and small letters.

• Capitalization of the headings for units within sections should follow the same conventions as capitalization in a standard English sentence. The same applies to captions for figures and tables.

• Headings should contain no font specifications other than for the mathematical expressions in them.

• All headings in the text should be generated either by sectioning commands or by the standard headings for a theorem-like environment or for a figure or table. Headings for sectional units are generated by the following commands.

AMS-L^{A}TEX:  AMS-TEX:

  \part{}                    \specialhead . . . \endspecialhead
  \section{}                 \head . . . \endhead
  \subsection{}              \subhead . . . \endsubhead
  \subsubsection{}           \subsubhead . . . \endsubsubhead

  \paragraph{}

2Exceptions to this and the preceding rule may be made for local redefinitions of parameters in a diagram package or for fine tuning of the behavior of \qed, as described in Section 2.

2. Theorem-like Environments

Different sorts of theorem-like environments are treated differently. The technicalities are important enough that we treat AMS-L^{A}TEX and AMS-TEX separately. But the principles are similar.

2.1. AMS-L^{A}TEX. The nyj style uses the macros in theorem.sty, by Frank Mittelbach. Building on these macros, nyj.sty gives three types of \newtheorem environment: The style of output is determined by the value given to \theoremstyle.

The default (\theoremstyle{plain}) produces output in which the statement is set in italic type. This is to be used for formal assertions of mathematical results, e.g., theorems, corollaries, lemmas and propositions.

For proper æsthetics, it is important to set any parentheses, brackets, or braces in the body of these statements in roman type.^{3} (This may be done with the control sequence \rom{}.) Similarly, numbers should be set in math mode.

We use \theoremstyle{remark} for remarks or other environments of formal discussion. It gives a heading in italics (e.g., “Remark 2.3”) followed by text in roman type.

The \theoremstyle{definition} is to be used for all other theorem-like environments that require \newtheorem declarations, e.g., definitions, notations, examples, conjectures, and problems. It gives a bold faced heading followed by roman type.

Environments for proofs do not need to be declared. There are two environments that may be used for proofs. The standard one is the pf environment. There is also a starred form, in which the heading can be altered. Thus,

\begin{pf*}{Proof of Theorem 1}

may be used if the statement of the theorem is separated from its proof.

The proof environments automatically place a Q.E.D. symbol ( ) after the proof. This can look awkward if the proof ends with a display or a list. In this case, insert a \qed by hand at the appropriate spot, and place a \renewcommand{\qed}{} immediately before the \end{pf} or \end{pf*}.

When declaring theorem-like environments, the theoremstyle used for a given \newtheorem statement will be the most recently declared theoremstyle, if there is one, or the default of plain if no theoremstyle has been declared. Thus, supposing no preceding theoremstyle has been defined,

\newtheorem{thm}{Theorem}[section]
\theoremstyle{definition}
\newtheorem{defn}[thm]{Definition}

defines a Theorem environment in plain theoremstyle, numbered by section, and a Definition environment in definition theoremstyle, numbered consecutively with the theorems.

We have placed a suggested set of newtheorem declarations in the template for AMS-L^{A}TEX papers.

3The same should be done in any other portion of the document not set in roman or bold type.

2.2. AMS-TEX. Here, formal assertions of mathematical results, such as theorems, corollaries, lemmas, and propositions, are to be set using \proclaim{} … \endproclaim, which uses italic type for the statement of results. Thus, for proper æsthetics, it is important to set any parentheses, brackets, or braces in the body of these statements in roman type.^{4} (This may be done with the control sequence \rom{}.) Similarly, numbers should be set in math mode.

The environment given by \remark{} … \endremark is for remarks or other environments of formal discussion, while \demo{} … \enddemo is used for proofs. The author should insert a \qed at the end of the proof (but before the \enddemo).

All other theorem-like environments, e.g., definitions, notations, examples, conjectures, and problems, are to be set using \definition{} … \enddefinition.

3. Commutative Diagrams, Tables, and Graphics

We shall permit a limited number of diagram packages to be used in our papers, all of them available in the archive, here. Currently, we shall permit the use of the packages of Paul Burchard, Michael Barr, and Kris Rose (xypic). The latter requires the installation of its own set of fonts, also available in our archive.

We shall also permit the inclusion, via dvipaste, of diagrams made using lamstex. Once again, the fonts are provided in the archive.

We shall also permit the inclusion of encapsulated postscript files, via epsf.tex, by Tomas Rokicki and Donald Knuth, which is also available in the archive. Large postscript files cannot, of course, be submitted via email. Please consult with the appropriate editor regarding submission of such files.

Large diagrams or graphics can cause trouble with page breaks, and should be included as a floating figure. It should be inserted at the top of a page, and the caption, if any, should be below the figure. This is accomplished as follows.

AMS-L^{A}TEX:  AMS-TEX:

      \begin{figure}[t]               \topinsert

⟨code for figure⟩  ⟨code for figure⟩
\caption{⟨caption text⟩}           \botcaption{⟨caption label⟩}
\end{figure}  ⟨caption text⟩

                                      \endcaption
                                      \endinsert

Similarly, large tables should be placed in floats set to print at the top of a page. In this case, the caption should be at the top of the float:

AMS-L^{A}TEX:  AMS-TEX:

      \begin{table}[t]                \topinsert

\caption{⟨caption text⟩}           \topcaption{⟨caption label⟩}
⟨code for table⟩  ⟨caption text⟩

      \end{table}                     \endcaption

⟨code for table⟩

                                      \endinsert

4The same should be done in any other portion of the document not set in roman or bold type.

4. Bibliographic References

Authors’ names should be given with first names before last. You may use initials in place of first names.

The title of the work referenced should be set in italic type. We adopt the convention that all major words in book titles are to be capitalized. For the title of an article, the rules of capitalization are the same as for a sentence of ordinary text. Any part or volume number associated with the title should be set in roman type, as should the edition of a book, if referenced.

In referencing a journal article, the volume number should be set in bold type and the year should be enclosed in parentheses. Thus, the entry for a paper in a journal should look like the entry for [Whi49] in the bibliography below. The entries for most books should look like those for [ZS58] or [Spi90]: The author, title, volume, edition, etc., should be followed by the name and location of the publisher and the date of publication.

If a book is part of a series of monographs, the series and number should be given, as in the entry for [Cha83] below.

Papers contained in books add an extra layer of complication. The title of the paper, in italics, is followed by the title of the book in roman type. The book title should be capitalized according to the convention for books. If the book is a conference proceedings, then the location and date of the conference should follow the book title in parentheses. The book’s editors should also be listed in parentheses. This is followed by the publication data for the book as described above, followed by the pagination for the article. See the citation below for [Wei87].

You may use numerical tags for the reference works, or mnemonic tags such as those used here. Citations of these references within the body of the text are generated by \cite{} commands in both AMS-L^{A}TEX and AMS-TEX.

In AMS-TEX, the bibliography is generated by the \Refs … \endRefs environment, as described in detail in [Spi90, Appendix C].

In AMS-L^{A}TEX, the bibliography is generated by the \bibitem{} entries in the thebibliography environment, as described in [Lam86, Section 4.3]. The key for a \cite call is the argument to \bibitem.

It can be convenient, when using AMS-L^{A}TEX, to keep bibliography databases from which a thebibliography environment can be automatically generated by the auxiliary program BibTEX. BibTEX generates the \bibitem entries from the \cite commands in the body of the paper, where the arguments to \cite are keys placed in bibliographic databases (.bib files) listed in your paper. This procedure has the advantage that the \bibitem entries are formatted according to a \bibliographystyle listed in your paper, which can be set to produce the formatting appropriate to the journal of your choice.

In particular, we have developed two bibliographystyle files for the New York Journal, nyjalpha.bst and nyjplain.bst, which generate mnemonic tags and numerical tags, respectively, in the style specified here. (The author must still exercise care in choosing correct capitalization for titles and correct abbreviations for journal names, etc.) A detailed description of the use of BibTEX can be found in [Lam86, Appendix B].

Authors are not compelled to use the default mnemonic labels produced by nyjalpha.bst. It should be noted that the order of the \bibitems produced by BibTEX is obtained from the mnemonic tags, so you may want to rearrange the entries as well.

The output from running BibTEX is written to a separate .bbl file. When submitting your paper, you should put the contents of the .bbl file into the paper itself, in place of the \bibliography and \bibliographystyle commands used to generate it.

Here is the code, as a .bib entry for use with nyjalpha.bst, that will generate the bibliographic reference for [CS79], below.

@inproceedings{pseudo-free,
author = "S. E. Cappell and J. L. Shaneson",
title = "Pseudo-free actions, \rom{I}",
booktitle = "Proc. Pure Math. \AA rhus 1978",
confdata = "\AA rhus, Denmark, 1978",
editor = "J. L. Dupont and I. H. Madsen",
series = "Lect. Notes in Math.", number = 763,
year = 1979, pages = "395--447",
publisher = "Springer-Verlag", address = "Berlin"}

The initial string “pseudo-free” is the argument to be used for the \cite command. The tag produced by BibTEX depends on the authors and title. The confdata field is unique to the nyj styles, and is used for the location and date of a conference. The use of the series and number fields in @inproceedings entries is not documented in [Lam86], as the original styles did not permit it.

In AMS-TEX, an equivalent entry would be produced by

\ref \key CS79 \by S. E. Cappell and J. L. Shaneson
\paper Pseudo-free actions, \rom{I}
\inbook Proc. Pure Math. \AA rhus 1978 (\AA rhus, Denmark, 1978)
\eds J. L. Dupont and I. H. Madsen
\bookinfo Lect. Notes in Math., no. 763
\publ Springer-Verlag \publaddr Berlin
\yr 1979 \pages 395--447 \endref

within the \Refs environment. Note that the tag is given by the \key command, and is chosen by the author, rather than generated by the program.

References

[AUG90] AMS-L^{A}TEXversion 1.1 user’s guide, American Mathematical Society, Providence, R.I., 1990.

[CS79] S. E. Cappell and J. L. Shaneson, Pseudo-free actions, I, Proc. Pure Math. Århus 1978 (Århus, Denmark, 1978) (J. L. Dupont and I. H. Madsen, eds.), Lect. Notes in Math., no. 763, Springer-Verlag, Berlin, 1979, pp. 395–447.

[Cha83] T. A. Chapman, Controlled Simple Homotopy Theory, Lect. Notes in Math., no. 1009, Springer-Verlag, Berlin, 1983.

[Knu84] Donald E. Knuth, The TEXbook, Addison-Wesley, Reading, Massachusetts, 1984.

[Lam86] Leslie Lamport, Latex: a Document Preparation System, Addison-Wesley, Reading, Massachusetts, 1986.

[Spi90] Michael D. Spivak, The Joy of TEX, second ed., American Mathematical Society, Providence, R.I., 1990.

[Wei87] Shmuel Weinberger, Continuous versus discrete symmetry, Geometry and Topology (Athens, Georgia, 1985) (C. McCrory and T. Shifrin, eds.), Marcel Dekker, New York, 1987, pp. 319–324.

Style for New York Journal of Mathematics 7

[Whi49] J. H. C. Whitehead, Combinatorial homotopy theory, I, Bull. Amer. Math. Soc. 55 (1949), 213–245.

[ZS58] Oscar Zariski and Pierre Samuel, Commutative Algebra, vol. I, Van Nostrand, Princeton, N.J., 1958.

Department of Mathematics and Statistics, The University at Albany, State University of New York, Albany, NY 12222

e-mail: mark@sarah.albany.edu